OpenStudy (anonymous):

The problem says: Find all pair of integers such that when they are added they equal 100 and when they are multiplied they are at least 2000. This is what I have so far: x+y=100 xy=2000

OpenStudy (cwrw238):

the second part should be xy >= 2000

OpenStudy (anonymous):

I solved for x: x=100-y then i substitute x into the second equation

OpenStudy (mathstudent55):

The expression "at least" is a key word that an inequality is being used.

OpenStudy (mathstudent55):

Continue what you are doing, but use @cwrw238 's inequality to plug it into.

OpenStudy (anonymous):

ok.. let me try that.

OpenStudy (anonymous):

(100-y)y > 2000

OpenStudy (cwrw238):

ok except its greater than or equal to >=

OpenStudy (anonymous):

\[(100-y)y \ge 2000\]

OpenStudy (cwrw238):

right

OpenStudy (cwrw238):

now froget the > for time being and solve the quadratic equation

OpenStudy (anonymous):

100y-\[y^{2}\] \[\ge 2000\]

OpenStudy (anonymous):

ok let me try to solve it

OpenStudy (anonymous):

let me re-do \[100y-y ^{2} \ge 2000\] Then I get \[-y ^{2} \ge 2000-100\]

OpenStudy (cwrw238):

you need to solve y^2 - 100y + 2000 = 0

OpenStudy (cwrw238):

this wont factor so you'll need to use quadratic formula

OpenStudy (cwrw238):

are you ok with doing that?

OpenStudy (cwrw238):

i'll be back in 3-4 minutes

OpenStudy (anonymous):

\[x= \frac{ -(-100) \pm \sqrt{(-100)^{2}-4 (1)(2000)} }{ 2 }\] \[x \frac{ 100 \pm \sqrt{1000- 8000} }{2}\]

OpenStudy (anonymous):

\[x= \frac{ 100 \pm \sqrt{-7000} }{ 2 }\]

OpenStudy (cwrw238):

right you'll find that comes to 72.36 and 27.64

OpenStudy (anonymous):

ok did i miss a step because i dont get 72.36 and 27.64

OpenStudy (cwrw238):

100^2 = 10,000

OpenStudy (cwrw238):

error there

OpenStudy (anonymous):

oh i see! thank you so much. I was trying to factor it. Now i now why i wasnt getting it.

OpenStudy (cwrw238):

so we see one pair of integers already 72 and 18 = they add up to 100 and 28*72 = 2016

OpenStudy (anonymous):

right! yes, i see it now. and since it says at least 2000 its ok to get 2016 right

OpenStudy (anonymous):

once i go back to check the answer i get 100 and 2016

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

but thats not all the possible pairs eg 71 and 29 gives 100 and 2019

OpenStudy (cwrw238):

I'm not quite sure how to continue form here....

OpenStudy (anonymous):

thats more than enough. Thank you so much!

OpenStudy (cwrw238):

ok - no problem

OpenStudy (cwrw238):

I got the graph on my calculator pairs of numbers which add up to 100 which are along the heavy line fit the bill